twice a number increased by 5 is 35.find the number
Answer by nerdybill(5410)

Let x = the number
then "twice a number increased by 5 is 35" translate to:
2x + 5 = 35
.
Solving for 'x':
2x + 5 = 35
Subtract 5 from both sides:
2x + 5 - 5 = 35 - 5
2x = 30
Finally, divide both sides by 2:
2x/2 = 30/2
x = 15 (this is your number)